OpenStudy (ilovecake):

Evaluate the expression when m = 4. If someone helps I will give a medal please help :) 35 – 3m A. 1 B. 13 C. 23 D. 128

OpenStudy (ilovecake):

someone help

OpenStudy (ilovecake):

I will give medal and I will become fan if you help

OpenStudy (ilovecake):

someone please help :(

OpenStudy (anonymous):

i think it's B

OpenStudy (ilovecake):

you sure?

OpenStudy (anonymous):

b 35-3*4= 35-12=13

OpenStudy (ilovecake):

Ok thanks to both of you :)